# A General Theory of Static Response for Markov Jump Processes

### v1

@@ Abstract
We consider Markov jump processes on a graph described by a rate matrix that depends on various control parameters. We derive explicit expressions for the static responses of edge currents and steady-state probabilities. We show that they are constrained by the graph topology (i.e. the incidence matrix) by deriving response relations (i.e. linear constraints linking the different responses) and topology-dependent bounds. For unicyclic networks, all scaled current sensitivities are between zero and one and must sum to one. Applying these results to stochastic thermodynamics, we derive explicit expressions for the static response of fundamental currents (which carry the full dissipation) to fundamental thermodynamic forces (which drive the system away from equilibrium).
Introduction.—Nonequilibrium steady state (NESS) of Markov jump processes describe a plethora of phenomena [1] and understanding their response to external perturbations has crucial implications [2] across fields, such as biology [3, 4, 5, 6], nanoelectronics [7, 8], and deep learning [9]. When the Markov jump process is produced by thermal noise, near equilibrium, the response is simple and characterized by the dissipation-fluctuation relation (DFR) [10, 11]. But far-from-equilibrium or for non-thermal processes, the response is significantly more involved; see Refs. [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Recently, for the static response, exact results [24, 25] and tight bounds [26, 27, 28] have been derived assuming Arrhenius-like rates. Moreover, the bounds [24] only hold for local responses, i.e. when the perturbation and the observable are assigned to the same transition.
@@
Setup.—We consider a directed graph $\mathcal{G}$ with $N$ nodes and $N_{e}$ edges and a Markov jump process over the discrete set $\mathcal{S}$ of the $N$ states corresponding to the nodes. Then, the edges $e\in\mathcal{E}$ of $\mathcal{G}$ define possible transitions with the probability rates encoded in the rate matrix $\mathbb{W}/\tau$. In this description, the jump from $n$ to $m$ is the edge (arrow) $+e$ with the source $s(+e)=n$ and the tip $t(e)=m$. For the reverse transition, we have $-e$ with $s(-e)=m$ and $t(-e)=n$. Choosing $\tau=1$ the non-diagonal elements $W_{nm}=W_{e}\geq 0$ become the probabilities per unit of time assigned to the edges $e$. We assume that the matrix $\mathbb{W}$ is irreducible [30] and that all transitions are reversible, i.e. $W_{e}\neq 0$ only if $W_{-e}\neq 0$. With the property of diagonal elements $W_{ii}=-\sum_{j\neq i}W_{ij}$ the described system always exhibits a unique steady-state probability $\bm{\pi}=(\pi_{1},\dots,\pi_{N})^{\intercal}$ that satisfies
@@
with the normalization $\sum_{i=1}^{N}{\pi_{i}}=1$. We define the transition current along the edge $e$ as $j_{e}=W_{+e}\pi_{s(+e)}-W_{-e}\pi_{s(-e)}$. This definition has a matrix form $\bm{j}=\mathbb{\Gamma}\bm{\pi}$, where the matrix $\mathbb{\Gamma}$ has elements $\Gamma_{ei}=W_{+e}\delta_{is(+e)}-W_{-e}\delta_{is(-e)}$ with the Kronecker symbol $\delta$. In analogy to linear chemical reaction networks [31, 32, 33], the rate matrix can be decomposed as $\mathbb{W}=\mathbb{S}\mathbb{\Gamma}$, where $\mathbb{S}$ is the incidence matrix of the directed graph $\mathcal{G}$ with the elements $S_{ie}=\delta_{is(-e)}-\delta_{is(+e)}$.
We introduce parameterization $\mathbb{W}(\bm{p})$ using the vector $\bm{p}=(\dots,p,\dots)^{\intercal}$ made of $N_{p}$ control parameters $p\in\bm{p}$. Thus, the steady-state condition (1) can be written as
@@
where $\bm{j}(\bm{p})\equiv\bm{j}(\bm{p},\bm{\pi}(\bm{p}))$ depends on $\bm{p}$ both explicitly due to $\mathbb{\Gamma}(\bm{p})$ and implicitly due to $\bm{\pi}(\bm{p})$. The central objects of this work are the static responses of a quantity $q(\bm{p})$ (e.g. a probability $\bm{\pi}$ or a current $\bm{j}$) to perturbations of the parameters $p$, i.e. the elements of the vector $\nabla_{\bm{p}}q(\bm{p})=(\dots,d_{p}q,\dots)^{\intercal}$.
Homogeneous parameterization.—To provide a simple illustration of how a given parameterization can give rise to nontrivial relations amongst static responses, let us consider a parameterization $\bm{h}$ of the rate matrix $\mathbb{W}(\bm{h})$ such that:
@@
$\sum_{h\in\bm{h}}hd_{h}\bm{j}$ $=k\bm{j}\,.$ (5)
Equations 4 and 5 are known in metabolic control analysis as summation theorems; see [34, 35, 36, 37]. In that context, enzyme concentrations play the role of the homogeneous parameters.
Matrix approach to static response.— We now turn to arbitrary parameterizations. Our strategy is to use $\sum_{i=1}^{N}\pi_{i}=1$, and thus $d_{p}\pi_{N}=-\sum_{k=1}^{N-1}d_{p}\pi_{k}$, to arrive at $N-1$ independent equations for others $d_{p}\pi_{k}$ with $k\in\hat{\mathcal{S}}=\mathcal{S}\setminus\{N\}$. To solve this linear problem, we introduce the matrix $\mathbb{K}=\hat{\mathbb{S}}\hat{\mathbb{\Gamma}}$, where $\hat{\mathbb{S}}=[S_{ke}]_{\{k,e\}}$ and $\hat{\mathbb{\Gamma}}=[\Gamma_{ek}-\Gamma_{eN}]_{\{e,k\}}$ are reduced matrices with $k\in\hat{\mathcal{S}}$ and $e\in\mathcal{E}$. In Appendix A, we prove that the matrix $\mathbb{K}$ is invertible [see Eq. 53], and derive the following expressions for the probability and current response matrices:
$\mathbb{R}^{\pi}$ $\equiv[d_{p}\pi_{i}]_{\{i,p\}}=-\begin{pmatrix}\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J}\\ -\sum_{k}(\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J})^{\intercal}_{k}\end{pmatrix}\,,$ (6c)
$\mathbb{R}^{j}$ $\equiv[d_{p}j_{e}]_{\{e,p\}}=\mathbb{P}\mathbb{J}\,.$ (6d)
Here, $\sum_{k}(\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J})^{\intercal}_{k}$ denotes the row that is the sum of all rows of the matrix $\mathbb{K}^{-1}\hat{S}\mathbb{J}$; $\{i,p\}$ (resp. $\{e,p\}$) denote the sets of indexes $i\in\mathcal{S}$ (resp. $e\in\mathcal{E}$) for rows and $p\in\bm{p}$ for columns; $\mathbb{J}\equiv[\partial_{p}j_{e}]_{\{e,p\}}$ is the steady state Jacobian
@@
$\mathbb{P}$ $\equiv\Big{[}\delta_{ee^{\prime}}-\sum_{x,x^{\prime}\in\hat{\mathcal{S}}}\hat{\Gamma}_{ex}(\mathbb{K}^{-1})_{xx^{\prime}}S_{x^{\prime}e^{\prime}}\Big{]}_{\{e,e^{\prime}\}}\,,$
$\mathbb{P}$ $=\mathbb{I}-\hat{\mathbb{\Gamma}}(\hat{\mathbb{S}}\hat{\mathbb{\Gamma}})^{-1}\hat{\mathbb{S}}\,,$ (8)
@@
where $\mathbb{C}=(\dots,\bm{c}_{\gamma},\dots)$ is the matrix of the fundamental cycles $\bm{c}^{\gamma}$ defined as the right null vectors of the incidence matrix $\mathbb{S}\bm{c}^{\gamma}=\bm{0}$ (note that $\mathbb{S}$ and $\hat{\mathbb{S}}$ share the same $N_{c}$ cycles). A General Theory of Static Response for Markov Jump Processes is crucial in what follows. In fact, the elements of the matrices $\mathbb{R}^{\pi}$ and $\mathbb{R}^{j}$ cover the responses of all edges to the perturbation of arbitrary parameters. The matrix structures on the r.h.s. of A General Theory of Static Response for Markov Jump Processes encode the information on how those responses are connected to each other.
@@
Response relations.—For a vector $\bm{p}$ such that the matrix $\mathbb{J}$ in Eq. 7 is full row rank ($\text{rk}\mathbb{J}=N_{e}$, i.e. $N_{p}\geq N_{e}$), we can always find the right invertible matrix $\mathbb{J}^{+}$ with $\mathbb{J}\mathbb{J}^{+}=\mathbb{I}$:
$\mathbb{J}^{+}=\mathbb{J}^{\intercal}(\mathbb{J}\mathbb{J}^{\intercal})^{-1}\,.$ (10)
@@
Multiplying both sides of Eqs. 6c and 6d on the vector $\mathbb{J}^{+}\bm{j}$, we arrive at $\mathbb{R}^{\pi}\mathbb{J}^{+}\bm{j}=\bm{0}$ and $\mathbb{R}^{j}\mathbb{J}^{+}\bm{j}=\mathbb{P}\bm{j}=\mathbb{I}\bm{j}$. In coordinate form, these relations give rise to the SRRs:
$\sum_{p\in\bm{p}}\phi_{p}d_{p}\bm{\pi}$ $=0\,,\quad\sum_{p\in\bm{p}}\phi_{p}d_{p}\ln\bm{j}=\bm{1}\,,$ (11)
where $d_{p}\ln\bm{j}=(\dots,j_{e}^{-1}d_{p}j_{e},\dots)^{\intercal}$ is the sensitivity vector, $\bm{1}=(\dots,1,\dots)^{\intercal}$, and $\phi_{p}$ are elements of the vector
$\bm{\phi}$ $=\mathbb{J}^{+}\cdot\bm{j}\,.$ (12)
@@
Equations (11) generalize Eqs. 4 and 5 beyond homogeneous parameters. The structure of Eq. 11 is universal, and the parameterization as well as the topology of the graph are encoded only in the vector Eq. 12, which can be calculated explicitly using Eq. 7. The SRR for the current in Eq. 11 constrains the extremum condition for the edge flux $j_{e}$ with respect to the vector $\bm{p}$: the condition $d_{p}j_{e}=0$ for $p\in\bm{p}$ is satisfied only for $j_{e}=0$ on the r.h.s. of Eq. 11. We stress that many physically motivated parameterisations satisfy Eq. 10.
@@
Edge perturbations.—We now restrict our theory to systems with independent parameters at every edge. In this case, every element $p_{e}$ of the vector $\bm{p}$ is assigned to its own edge $e$, which implies that the matrix $\mathbb{J}(\bm{p})=\text{diag}(\dots,\partial_{p_{e}}j_{e},\dots)$ is diagonal and $\mathbb{J}^{+}=\mathbb{J}^{-1}$. Using Eq. 12, we find
$\phi_{e}=(\partial_{p_{e}}j_{e})^{-1}j_{e}\,.$ (13)
With Eq. 13 at hand, we can find analytic expressions for Eq. 11. Moreover, since the matrix $\mathbb{J}$ is diagonal, we can rewrite Eq. 6d as
$P_{ee^{\prime}}=(\phi_{e^{\prime}}/j_{e^{\prime}})d_{p_{e^{\prime}}}j_{e}\,.$ (14)
Equation 14 shows that the elements $P_{ee^{\prime}}$ are scaled responses. In addition, we define the scaled sensitivity as $\phi_{e^{\prime}}d_{p_{e^{\prime}}}\ln j_{e}=(j_{e^{\prime}}/j_{e})P_{ee^{\prime}}$. Thus, we can characterize the responses and sensitivities from the analysis of the projection matrix $\mathbb{P}$.
@@
For edge parameterization, we call the responses $d_{p_{e^{\prime}}}j_{e}$ and $P_{ee^{\prime}}$ nonlocal if the perturbation edge $e^{\prime}$ does not coincide with the observation edge $e$. In Appendix C we prove that the diagonal elements of $\mathbb{P}$ are bounded:
@@
which implies the bounds for the local sensitivities [$e^{\prime}=e$ in Eq. 14]:
$0\leq\phi_{e}d_{p_{e}}\ln j_{e}\leq 1\,.$ (16)
Equation 16 holds for an arbitrary edge parameterization that extends the bounds found in Ref. [24]. In fact, the bounds in Ref. [24] assume an Arrhenius-like form of the edge perturbations (special form of $\bm{\phi}$). We will discuss them as in the application section, see Fig. 1.
@@
Using the property of idempotent matrices $\text{tr}~{}\mathbb{P}=\text{rk}~{}\mathbb{P}=N_{c}$, we derive the CRRs:
$\sum_{e}\phi_{e}d_{p_{e}}\ln j_{e}=N_{c}\,,$ (17)
@@
where the l.h.s. is the sum of local sensitivities which are bounded by Eq. 16. Therefore, Eqs. 16 and 17 show that if exactly $N_{c}$ local sensitives are saturated, then the other ones must be zero. We will illustrate Eq. 17 with an example in the application to physics section.
@@
Figure 1: (a): Sketch of unimolecular network. (b): The heights of the purple, blue, green, yellow bins correspond to the scaled responses of edge $1$ to perturbation of $F_{1},\dots,F_{4}$, with $W_{\pm e}=\exp(v_{e}\pm F_{e})$ and $\phi_{e}=\tau_{e}/j_{e}$. The responses calculated for $100$ rate-matrices with random $W_{\pm e}$ homogeneously distributed in $0\leq W_{\pm e}\leq 100$.

### v2

# General Theory of Static Response for Markov Jump Processes

@@ Abstract
We consider Markov jump processes on a graph described by a rate matrix that depends on various control parameters. We derive explicit expressions for the static responses of edge currents and steady-state probabilities. We show that they are constrained by the graph topology (i.e. the incidence matrix) by deriving response relations (i.e. linear constraints linking the different responses) and topology-dependent bounds. For unicyclic networks, all scaled current responses are between zero and one and must sum to one. Applying these results to stochastic thermodynamics, we derive explicit expressions for the static response of fundamental currents (which carry the full dissipation) to fundamental thermodynamic forces (which drive the system away from equilibrium).
Introduction.—Nonequilibrium steady state (NESS) of Markov jump processes describe a plethora of phenomena [1] and understanding their response to external perturbations has crucial implications [2] across fields, such as biology [3, 4, 5, 6], nanoelectronics [7, 8], and deep learning [9]. When the Markov jump process is produced by thermal noise, near equilibrium, the response is simple and characterized by the dissipation-fluctuation relation (DFR) [10, 11]. But far-from-equilibrium or for non-thermal processes, the response is significantly more involved; see Refs. [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Recently, for the static response, exact results [25, 26, 27] and tight bounds [28, 29] have been derived assuming Arrhenius-like rates. Moreover, the bounds [25] only hold for local responses, i.e. when the perturbation and the observable are assigned to the same transition.
@@
Setup.—We consider a directed graph $\mathcal{G}$ with $N$ nodes and $N_{e}$ edges and a Markov jump process over the discrete set $\mathcal{S}$ of the $N$ states corresponding to the nodes. Then, the edges $e\in\mathcal{E}$ of $\mathcal{G}$ define possible transitions with the probability rates encoded in the rate matrix $\mathbb{W}/\tau$. In this description, the jump from $n$ to $m$ is the edge (arrow) $+e$ with the source $s(+e)=n$ and the tip $t(+e)=m$. For the reverse transition, we have $-e$ with $s(-e)=m$ and $t(-e)=n$. Choosing $\tau=1$ the non-diagonal elements $W_{nm}=W_{e}\geq 0$ become the probabilities per unit of time assigned to the edges $e$. We assume that the matrix $\mathbb{W}$ is irreducible [31] and that all transitions are reversible, i.e. $W_{e}\neq 0$ only if $W_{-e}\neq 0$. With the property of diagonal elements $W_{ii}=-\sum_{j\neq i}W_{ij}$ the described system always exhibits a unique steady-state probability $\bm{\pi}=(\pi_{1},\dots,\pi_{N})^{\intercal}$ that satisfies
@@
with the normalization $\sum_{i=1}^{N}{\pi_{i}}=1$. We define the transition current along the edge $e$ as $j_{e}\equiv W_{+e}\pi_{s(+e)}-W_{-e}\pi_{s(-e)}$. This definition has a matrix form $\bm{j}=\mathbb{\Gamma}\bm{\pi}$, where the matrix $\mathbb{\Gamma}$ has elements $\Gamma_{ei}\equiv W_{+e}\delta_{is(+e)}-W_{-e}\delta_{is(-e)}$ with the Kronecker symbol $\delta$. In analogy to linear chemical reaction networks [32, 33, 34], the rate matrix can be decomposed as $\mathbb{W}=\mathbb{S}\mathbb{\Gamma}$, where $\mathbb{S}$ is the incidence matrix of the directed graph $\mathcal{G}$ with the elements $S_{ie}\equiv\delta_{is(-e)}-\delta_{is(+e)}$.
We introduce a parameterization $\mathbb{W}(\bm{p})$ using the vector $\bm{p}=(\dots,p,\dots)^{\intercal}$ made of $N_{p}$ control parameters $p\in\bm{p}$. Thus, the steady-state condition (1) can be written as
@@
where $\bm{j}(\bm{p})\equiv\bm{j}(\bm{p},\bm{\pi}(\bm{p}))$ depends on $\bm{p}$ both explicitly due to $\mathbb{\Gamma}(\bm{p})$ and implicitly due to $\bm{\pi}(\bm{p})$. The central objects of this work are the static responses of a quantity $q(\bm{p})$ (e.g. a probability $\bm{\pi}$ or a current $\bm{j}$) to perturbations of the parameters $p$, i.e. the elements of the vector $\nabla_{\bm{p}}q(\bm{p})=(\dots,d_{p}q,\dots)^{\intercal}$. Our goal will be to obtain explicit relations for these responses enabling us to find relations among them. Many recent works considered edge perturbations where each element of $\bm{p}$ acts on a rate associated with a given edge [25, 35, 36, 27, 26] . But often, perturbing physical quantities implies acting on rates associated with many edges. In this letter, after an example, we go from generic perturbations to more specific ones, illustrating our findings at every stage using a physical model.
Homogeneous parameterization.—We first provide a simple illustration of how a given parameterization can give rise to nontrivial relations amongst static responses, let us consider the parameterization $\bm{h}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\dots,h_{p},\dots)^{\intercal}}$ of $\mathbb{W}(\bm{h})$ such that:
@@
$\sum_{p}h_{p}\frac{d\bm{j}}{dh_{p}}$ $=k\bm{j}\,.$ (5)
Equations 4 and 5 are known in metabolic control analysis as summation theorems [37, 38, 39, 40]. In that context, enzyme concentrations play the role of the homogeneous parameters.
Matrix approach to static response.— We now turn to arbitrary parameterizations. Our strategy is to use $\sum_{i=1}^{N}\pi_{i}=1$, and thus $d_{p}\pi_{N}=-\sum_{k=1}^{N-1}d_{p}\pi_{k}$, to arrive at $N-1$ independent equations for others $d_{p}\pi_{k}$ with $k\in\hat{\mathcal{S}}\equiv\mathcal{S}\setminus\{N\}$. To solve this linear problem, we introduce the matrix $\mathbb{K}\equiv\hat{\mathbb{S}}\hat{\mathbb{\Gamma}}$, where $\hat{\mathbb{S}}\equiv[S_{ke}]_{\{k,e\}}$ and $\hat{\mathbb{\Gamma}}\equiv[\Gamma_{ek}-\Gamma_{eN}]_{\{e,k\}}$ are reduced matrices with $k\in\hat{\mathcal{S}}$ and $e\in\mathcal{E}$. In Sec. A of [41] we prove that the matrix $\mathbb{K}$ is invertible [see Eq. A5], and that the probability and current response matrices read:
$\mathbb{R}^{\pi}$ $\equiv[d_{p}\pi_{i}]_{\{i,p\}}=-\begin{pmatrix}\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J}\\ -\sum_{k}(\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J})^{\intercal}_{k}\end{pmatrix}\,,$ (6a)
$\mathbb{R}^{j}$ $\equiv[d_{p}j_{e}]_{\{e,p\}}=\mathbb{P}\mathbb{J}\,.$ (6b)
Here, $\sum_{k}(\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J})^{\intercal}_{k}$ denotes the row that is the sum of all rows of the matrix $\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J}$; $\{i,p\}$ (resp. $\{e,p\}$) denote the sets of indexes $i\in\mathcal{S}$ (resp. $e\in\mathcal{E}$) for rows and $p\in\bm{p}$ for columns; $\mathbb{J}\equiv[\partial_{p}j_{e}]_{\{e,p\}}$ is the steady state Jacobian
@@
$\mathbb{P}$ $\equiv\Big{[}\delta_{ee^{\prime}}-\sum_{x,x^{\prime}\in\hat{\mathcal{S}}}\hat{\Gamma}_{ex}(\mathbb{K}^{-1})_{xx^{\prime}}S_{x^{\prime}e^{\prime}}\Big{]}_{\{e,e^{\prime}\}}=\mathbb{I}-\hat{\mathbb{\Gamma}}(\hat{\mathbb{S}}\hat{\mathbb{\Gamma}})^{-1}\hat{\mathbb{S}}\,,$ (8)
@@
where $\mathbb{C}=(\dots,\bm{c}_{\gamma},\dots)$ is the matrix of the fundamental cycles $\bm{c}^{\gamma}$ defined as the right null vectors of the incidence matrix $\mathbb{S}\bm{c}^{\gamma}=\bm{0}$ ($\mathbb{S}$ and $\hat{\mathbb{S}}$ share the same $N_{c}$ cycles). Equation 6 is crucial in what follows. It contains explicit expressions for the responses of all edges to arbitrary perturbations and contains information on how they are related to each other.
@@
Response relations.—For a vector $\bm{p}$ such that the matrix $\mathbb{J}$ in Eq. 7 is full row rank ($\text{rk}\mathbb{J}=N_{e}$, i.e. $N_{p}\geq N_{e}$), we can always find a right invertible matrix $\mathbb{J}^{+}$ such that $\mathbb{J}\mathbb{J}^{+}=\mathbb{I}$:
$\mathbb{J}^{+}\equiv\mathbb{J}^{\intercal}(\mathbb{J}\mathbb{J}^{\intercal})^{-1}\,.$ (10)
@@
Multiplying both sides of Eqs. 6a and 6b by the vector $\mathbb{J}^{+}\bm{j}$, we arrive at $\mathbb{R}^{\pi}\mathbb{J}^{+}\bm{j}=\bm{0}$ and $\mathbb{R}^{j}\mathbb{J}^{+}\bm{j}=\mathbb{P}\bm{j}=\mathbb{I}\bm{j}$. In coordinate form, these relations give rise to the SRRs:
$\sum_{p\in\bm{p}}\phi_{p}d_{p}\pi_{i}$ $=0\,,\quad\sum_{p\in\bm{p}}\phi_{p}d_{p}j_{e}=j_{e}\,,$ (11)
where $\phi_{p}$ are elements of the vector
$\bm{\phi}$ $\equiv\mathbb{J}^{+}\cdot\bm{j}\,.$ (12)
@@
Equations (11) generalize Eqs. 4 and 5 beyond homogeneous parameters with the only difference that the $h_{p}$’s are now replaced by the coefficients $\phi_{p}$. The structure of Eq. 11 is universal, and the parameterization as well as the topology of the graph are encoded only in the vector Eq. 12, which can be easily calculated explicitly using Eq. 7. The SRR for the current in Eq. 11 constrains the response of the edge fluxes. It implies that all flux responses of a given edge $e$ vanish ($d_{p}j_{e}=0$ for all $p\in\bm{p}$) only when that edge is at equilibrium, i.e. when $j_{e}=0$.
Figure 1: (a) Double QDs u and d coupled with three reservoirs (purple, blue, green). The reservoirs have different temperatures $T_{i}$ ($\beta_{i}=1/T_{i}$) and chemical potentials $\mu_{i}$, where $i=1,2,3$. (b) Graph representation: Four states $00$, $01$, $10$, and $11$ have energies 00, $\epsilon_{u}$, $\epsilon_{d}$, and $\epsilon_{d}+\epsilon_{u}+U_{C}$, respectively, where $U_{C}$ is the Coulomb repulsion energy arising when the two dots are filled. The colored arrows show the transitions assigned to the corresponding reservoir. The transition rates read $W_{\pm e}=\Gamma_{e}(1+\exp(\pm\Psi_{e}))^{-1}$, where $\Gamma_{e}$ are the tunneling rates and $\Psi_{e}$ are the edge parameters (the potential assigned to the transition $+e$): $\Psi_{1}=\beta_{1}(\epsilon_{u}-\mu_{1})$, $\Psi_{2}=\beta_{2}(\epsilon_{d}-\mu_{2})$, $\Psi_{3}=\beta_{2}(\epsilon_{d}+U_{C}-\mu_{2})$, $\Psi_{4}=\beta_{1}(\epsilon_{u}+U_{C}-\mu_{1})$, $\Psi_{5}=\beta_{3}(\epsilon_{d}-\mu_{3})$ and $\Psi_{6}=\beta_{3}(\epsilon_{d}+U_{C}-\mu_{3})$.
Physical example: We consider the double Quantum Dots (QDs) model shown in Fig. 1[43, 44, 45, 46]. Each QD consists of a single electronic level with energy $\epsilon_{u}$, resp. $\epsilon_{d}$, that can be filled or empty due to electron exchanges with the reservoirs. Electrons cannot be transferred between the two QDs, but when the two QDs are occupied, they interact with each other via a Coulomb repulsion energy $U_{C}$. The four manybody states of the system and their corresponding energy are shown in Fig. 1b. In this case, the general explicit expressions for responses, Eq. 6, hold for any parameterization. But to satisfy the SRRs, we must have at least $6$ independent model parameters ($N_{p}\geq N_{e}=6$ and $\mathbb{J}$ must be full row rank): The set $\{\epsilon_{u},\epsilon_{d},U_{C}\}$ is not large enough [Eq. (B2) in [41]], the set $\{\epsilon_{u},\epsilon_{d},U_{C},\mu_{1},\mu_{2},\mu_{3}\}$ is not independent as $\det\mathbb{J}=0$ [Eq. (B3) in [41]], but the set $\{\epsilon_{u},\epsilon_{d},U_{C},\beta_{1},\beta_{2},\beta_{3}\}$ would work as $\det\mathbb{J}\neq 0$ [Eq. (B4) in [41]].
@@
Independent edge perturbations.—We now restrict our theory to systems with independent parameters at every edge, namely $\text{rk}~{}\mathbb{J}=N_{p}=N_{e}$. In this case, every element $p_{e}$ of the vector $\bm{p}$ is assigned to its own edge $e$, which implies that the matrix $\mathbb{J}(\bm{p})=\text{diag}(\dots,\partial_{p_{e}}j_{e},\dots)$ is diagonal and $\mathbb{J}^{+}=\mathbb{J}^{-1}$. Using Eq. 12, the coefficients $\phi_{e}$ take the explicit form
$\phi_{e}=\Big{(}\frac{\partial j_{e}}{\partial p_{e}}\Big{)}^{-1}j_{e}\,.$ (13)
Since the matrix $\mathbb{J}$ is diagonal, we can rewrite Eq. 6b as
$P_{ee^{\prime}}=\Big{(}\frac{\partial j_{e^{\prime}}}{\partial p_{e^{\prime}}}\Big{)}^{-1}\frac{dj_{e}}{dp_{e^{\prime}}}\,.$ (14)
This shows that the elements $P_{ee^{\prime}}$ are scaled responses where the scaling factor $\partial_{p_{e^{\prime}}}j_{e^{\prime}}$ is controlled by the explicit dependence $W_{\pm e}(\bm{p})$ and can be interpreted as the instantaneous response, as the edge probabilities $\pi_{s(\pm e^{\prime})}$ had no time to change. This means that $P_{ee^{\prime}}$ can be seen as the ratio between the complete and instantaneous response.
Physical example: For the QDs of Fig. 1, edge perturbation can be realized using the set $\bm{\Gamma}=(\dots,\Gamma_{e},\dots)^{\intercal}$. The set $\{\epsilon_{u},\epsilon_{d},U_{C},\beta_{1},\beta_{2},\beta_{3}\}$ also works if controlled in such a way that the edge parameters $\bm{\Psi}=(\dots,\Psi_{e},\dots)^{\intercal}$ are changed independently, see Sec. B of [41].
@@
For edge parameterization, we call the responses $d_{p_{e^{\prime}}}j_{e}$ and $P_{ee^{\prime}}$ local (resp. nonlocal) if the perturbation edge $e^{\prime}$ does (resp. does not) coincide with the observation edge $e$. In Sec. C of [41] we prove that the diagonal elements of $\mathbb{P}$ are bounded,
@@
which implies the bounds for the local scaled responses
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0\leq\Big{(}\frac{\partial j_{e}}{\partial p_{e}}\Big{)}^{-1}\frac{dj_{e}}{dp_{e}}\leq 1}\,.$ (16)
This result extends a previous bounds obtained in Ref. [25] for specific parameterisations and perturbations of the rates.
@@
Using the property of idempotent matrices $\text{tr}~{}\mathbb{P}=\text{rk}~{}\mathbb{P}=N_{c}$, we can further derive the following CRRs:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{e=1}^{N_{e}}\Big{(}\frac{\partial j_{e}}{\partial p_{e}}\Big{)}^{-1}\frac{dj_{e}}{dp_{e}}=N_{c}\,.}$ (17)
@@
Since the l.h.s. is the sum of local scaled responses which are bounded by Eq. 16, Eq. 17 show that if exactly $N_{c}$ local sensitives are saturated, then the other ones must be zero. This will be illustrated later in Fig. 3.
@@
Figure 2: (a) Unicyclic network obtained by removing the third reservoir from Fig. 1. (b) The heights of the purple, blue, green, yellow bins correspond to the scaled responses of $j_{1}$ to perturbation of $\Psi_{1},\dots,\Psi_{4}$ for $\Gamma_{e}$, $\beta_{1}$, $\beta_{2}$ randomly and homogeneously distributed in $0\leq\Gamma_{e}\leq 100$ and $0\leq\beta_{1},\beta_{2}\leq 2$, with $U_{C}=1$, $\mu_{1}=\mu_{2}=1$. In (c), resp. (d), $(T_{1}d_{\epsilon_{u}}j_{1}+|\partial_{\Psi_{1}}j_{1}|)/(|\partial_{\Psi_{1}}j_{1}|+|\partial_{\Psi_{4}}j_{4}|)\,$, resp. $(T_{2}d_{\epsilon_{d}}j_{1}+|\partial_{\Psi_{3}}j_{3}|)/(|\partial_{\Psi_{2}}j_{2}|+|\partial_{\Psi_{3}}j_{3}|)$, bounded between 00 and $1$.